f(x)=2x^2+x and g(x)=x-2 find f(g(x))
1
Expert's answer
2013-05-22T10:02:52-0400
To find composition of functions f(g(x)) we should substitute g(x)=x-2 instead of x into the formula for f(x):
f(g(x))=2(x-2)^2+(x-2)=2(x^2-4x+4)+x-2=2x^2-8x+8+x-2=2x^2-7x+6

To get more info about composition of functions watch our video on this stuff: